In a family there are two cars. In a given week, the first car gets an average of 15 miles per gallon, and the second car gets 35 miles per gallon. The two cars combined drive a total of 1575 miles in that week, for a total gas consumption of 65 gallons. How many gallons were consumed by each of the two cars that week?
Answer by mananth(16949)

In a family there are two cars. In a given week, the first car gets an average of 15 miles per gallon, and the second car gets 35 miles per gallon. The two cars combined drive a total of 1575 miles in that week, for a total gas consumption of 65 gallons. How many gallons were consumed by each of the two ?
..
let one car drive x miles in the week
the other car will drive 1575-x miles in the week.
..
first car drove 15x miles
second drove (1575-x)miles
..
x/15+(1575-x)/35 = 65
35x+15(1575-x) / 15*35= 65
35x+23625-15x=15*35*65
20x=34125-23625
20x=10500
x=525 miles driven by first car
gallons consumed = 525/15= 35
..
1575-525= 1050 miles traveled by second car
1050/ 35
=30 gallons